Fluorite, a mineral of calcium, is a compound of the metal with fluorine. Analysis shows that a 2.60 g sample of fluorite contains 1.67 g calcium. Calculate the following. (a) mass of fluorine in the sample g (b) mass fraction of calcium and fluorine in fluorite calcium fluorine (c) mass percent of calcium and fluorine in fluorite calcium % fluorine %?
Answer #1

fluorite is CaF2.

mass of Ca = 1.67 g

mass of F = (2.60 - 1.67) = 0.93 g

(a) mass of fluorine in the sample = 0.93 g

(b) mass fraction of calcium = 1.67 / 2.60 = 0.642

and

mass fraction of fluorine = 0.93 / 2.60 = 0.358

(c) mass percent of calcium = 1.67 * 100 / 2.60 = 64.2 %

and

mass percent of fluorine = 0.93 * 100 / 2.60 = 35.8 %